# KiDS-Legacy: angular galaxy clustering from deep surveys with complex selection effects

### v1

@@ Appendix C SOM+HC with all the systematics from KiDS-Legacy
Our fiducial choice of five systematics is based on the consideration that some systematics are strongly correlated with each other (such as extinction in different bands), so introducing them will not add information to the SOM. In addition, some systematics have no selection effect in the galaxy field. Therefore, including them in the training vector will not improve performance, but will increase the time and computational resources required (especially in the validation procedure). In this section, we validate our fiducial choice of systematics described in Section 3 by training the SOM with all the available systematics in the KiDS-Legacy catalogue. This increases the number of systematics from 5 to 16 and also increases the number of training epochs required for the SOM to converge.
We group the galaxy into 600 clusters as the fiducial setup and calculate Spearman’s correlation coefficient between median systematics and galaxy contrast in each hierarchical cluster. Figure 23 shows the correlation coefficient matrix. The first row of the correlation matrix shows the correlation between galaxy number contrast and each systematics, indicating the selection effect captured by SOM+HC. The extinction in different bands is fully correlated because it is calculated by scaling the reddening template given by Schlafly & Finkbeiner (2011) according to the band, so we only include $r$-band here as the fiducial case. PSF size has the highest negative correlation, which is in agreement with the fiducial run. Level and GAIA star number density are both slightly correlated with galaxy number contrast. We do not include Background in our fiducial run because it correlates weakly with galaxy number contrast, while it correlates strongly with GAIA star number density, so including it does not add much information.
To further justify our choice of systematics, we measure the 2PCF corrected by the OR weight generated from all the systematics. Figure 24 shows the UR (pink), fiducial OR (blue) and full systematics OR (orange) 2PCFs. The $\chi_{\mathrm{d}}^{2}$ value between the fiducial 2PCF and all difference is insignificant over the entire angular scale considered. Therefore, we conclude that the 5 systematics that we choose are representative of the whole set of systematics to recover the organised random.
@@ Appendix D Angular galaxy clustering for the KiDS-1000 bright sample
[…]
table_k1000b.tex
@@
Figure 26: The 2PCFs measured from the KiDS-1000 bright sample. The three panels are the auto-correlation $w(\theta)\,$of the whole sample, the first and the second tomographic bins, respectively. Blue and pink dots are the measurements corrected by recovered organised random and uniform random. The shaded regions are angular scales corresponding to physical scales smaller than $8h^{-1}\mathrm{Mpc}$ estimated at the mean redshift. The error bars are the standard deviations derived from the covariance matrix provided by the OneCovariance package. The orange curve is the OR-corrected 2PCF data points measured by Johnston et al. (2021). The black dashed curves in the second and third panels are the best-fit 2PCF from MCMC.
Figure 27: The posterior of $\Omega_{\mathrm{m}},b_{1},b_{2}$ fit by the 2PCF from KiDS-1000 bright samples. The deep-shaded and lightly-shaded contour is the 68% and 95% credible levels, respectively. The blue and pink contours correspond to the posterior from 2PCF corrected with OR and UR respectively.
Johnston et al. (2021) measured the 2PCF from the KiDS-1000 Bright sample (Bilicki et al. 2021) selected with a magnitude cut $r<20$ from the 4th KiDS data release. Its redshift distribution is reliably calibrated from the overlap with Galaxy And Mass Assembly (GAMA) spectroscopy using the neural network algorithm implemented in the ANNz2 software (see Figure 25 for the redshift distribution). The sample covers a sky area of 789 $\mathrm{deg}^{2}$ and has a number density of 0.36 $\mathrm{arcmin}^{-2}$.
In this section, we re-measure the 2PCF of the KiDS-Bright sample with the new SOM+HC implementation to check the consistency between our code and the pipeline used in Johnston et al. (2021). Following the fiducial setup (the “100A” setup) in Johnston et al. (2021), we recover the OR with a 100$\times$100 SOM trained on the same systematics ($r$-band detection threshold, PSF size and PSF shape) grouped into 100 hierarchical clusters. The OR weight map has a Nside=1024, which corresponds to an angular resolution of 3.4 arcmin, slightly larger than the fiducial setup in Johnston et al. (2021). The correlation function is then measured in 30 angular bins between 3 and 300 arcmin.
@@
After recovering the OR weight, we measure the correlation functions of the two tomographic bins of the bright sample, defined by selecting galaxies with ANNz-calibrated photo-$z$ with the cut $\{0.02,0.2,0.5\}$. The 2PCFs corrected by the uniform random and the recovered organised random are presented in Figure 26 with pink points and blue points, respectively. We notice that our measurements are fairly consistent with that from Johnston et al. (2021), shown as orange curves. The error bars are the standard deviation derived from the covariance matrix given by the OneCovariance code with the same redshift distribution and the best-fit $\Omega_{\mathrm{m}}$ and galaxy biases to be determined below. The shaded regions are angular scales corresponding to physical scales smaller than $8h^{-1}\mathrm{Mpc}$ estimated at the mean redshift. We then fit $\{\Omega_{\mathrm{m}},b_{1},b_{2}\}$ in the linear model with the 2PCFs in the linear scale. Note that we don’t know the parameters a priori to compute the covariance matrix, so we do an iterative parameter fit. We choose the initial value of $\Omega_{\mathrm{m}}\,$=0.33 and $b_{1}=1.1$, $b_{2}=1.25$ to compute the covariance matrix. The pilot galaxy bias values are taken from van Uitert et al. (2018), which apply to a GAMA-like subsample of KV-450 with a similar redshift distribution to the bright sample used here. We define a Gaussian likelihood and run an MCMC to obtain the best-fit parameters, and then update the covariance matrix with them. We then run the MCMC again with the updated covariance matrix to obtain the posterior.
@@
The posteriors of the UR and OR cases are shown as pink and blue contours in Figure 27. The mean value of the parameters in the converged MCMC chains and the $1-\sigma$ levels are summarized in Table 4. The theoretical $w(\theta)$ calculated from the best-fit parameters of the OR case is shown as dashed black curves in Figure 26. The reduced $\chi^{2}$ value between the OR 2PCF and the best-fit 2PCF is 1.06, corresponding to a PTE of 0.38, indicating a good fit between the model and the data. All three parameters are constrained. Notably, the matter density is constrained in agreement with previous cosmological probes. The galaxy biases are constrained to be close to 1 with a slightly increasing trend. However, we note to the reader that the bias parameters are highly degenerate with $\sigma_{8}$, which is fixed in our case. To break this degeneracy we need to introduce matter field tracers such as cosmic shear or use the halo model like in Dvornik et al. (2023). This is the content of the ongoing KiDS-Legacy $3\times 2$pt and $6\times 2$pt projects.
@@
Table 4: Summary of parameters fit from KiDS-1000 bright galaxy 2PCF. The best-fit values are those with maximum posterior and the errors are the 68.3% credible level.

### v2

@@ Appendix C SOM+HC with all the systematics from KiDS-Legacy
Our fiducial choice of five systematics is based on the consideration that some systematics are strongly correlated with each other (such as extinction in different bands), so introducing them will not add information to the SOM. In addition, some systematics have no selection effect in the galaxy field. Therefore, including them in the training vector will not improve performance, but will increase the time and computational resources required (especially in the validation procedure). In this section we validate our fiducial choice of systematics described in Section 3 by training the SOM with all the available systematics in the KiDS-Legacy catalogue. This increases the number of systematics from 5 to 16 and also increases the number of training epochs required for the SOM to converge.
We group the galaxy into 600 clusters as the fiducial set-up and calculate Spearman’s correlation coefficient between median systematics and galaxy contrast in each hierarchical cluster. Figure 23 shows the correlation coefficient matrix. The first row of the correlation matrix shows the correlation between galaxy number contrast and each systematics, indicating the selection effect captured by SOM+HC. The extinction in different bands is fully correlated because it is calculated by scaling the reddening template given by Schlafly & Finkbeiner (2011) according to the band, so we only include $r$-band here as the fiducial case. PSF size has the highest negative correlation, which is in agreement with the fiducial run. Level and GAIA star number density are both slightly correlated with galaxy number contrast. We do not include Background in our fiducial run because it correlates weakly with galaxy number contrast, while it correlates strongly with GAIA star number density, so including it does not add much information.
To further justify our choice of systematics, we measure the 2PCF corrected by the OR weight generated from all the systematics. Figure 24 shows the UR (pink), fiducial OR (blue) and full systematics OR (orange) 2PCFs. The $\chi_{\mathrm{d}}^{2}$ value between the fiducial 2PCF and all differences are insignificant over the entire angular scale considered. Therefore, we conclude that the 5 systematics that we choose are representative of the whole set of systematics to recover the organised random.
@@ Appendix D Angular galaxy clustering for the KiDS-1000 bright sample
Johnston et al. (2021) measured the 2PCF from the KiDS-1000 Bright sample (Bilicki et al. 2021) selected with a magnitude cut $r<20$ from the 4th KiDS data release. Its redshift distribution is reliably calibrated from the overlap with Galaxy And Mass Assembly (GAMA) spectroscopy using the neural network algorithm implemented in the ANNz2 software (see Fig. 25 for the redshift distribution). The sample covers a sky area of 789 $\mathrm{deg}^{2}$ and has a number density of 0.36 $\mathrm{arcmin}^{-2}$.
In this section we re-measure the 2PCF of the KiDS-Bright sample with the new SOM+HC implementation to check the consistency between our code and the pipeline used in Johnston et al. (2021). Following the fiducial set-up (the “100A” set-up) in Johnston et al. (2021), we recover the OR with a 100$\times$100 SOM trained on the same systematics ($r$-band detection threshold, PSF size and PSF shape) grouped into 100 hierarchical clusters. The OR weight map has a Nside=1024, which corresponds to an angular resolution of 3.4 arcmin, slightly larger than the fiducial set-up in Johnston et al. (2021). The correlation function is then measured in 30 angular bins between 3 and 300 arcmin.
@@
Figure 26: 2PCFs measured from the KiDS-1000 bright sample. The three panels are the auto-correlation $w(\theta)\,$of the whole sample, and the first and the second tomographic bins, respectively. The blue and pink dots are the measurements corrected by recovered organised random and uniform random. The shaded regions are angular scales corresponding to physical scales smaller than $8h^{-1}\mathrm{Mpc}$ estimated at the mean redshift. The error bars are the standard deviations derived from the covariance matrix provided by the OneCovariance package. The orange curve is the OR-corrected 2PCF data points measured by Johnston et al. (2021). The black dashed curves in the second and third panels are the best-fit 2PCF from MCMC.
@@
After recovering the OR weight, we measure the correlation functions of the two tomographic bins of the bright sample, defined by selecting galaxies with ANNz-calibrated photo-$z$ with the cut $\{0.02,0.2,0.5\}$. The 2PCFs corrected by the uniform random and the recovered organised random are presented in Fig. 26 with pink points and blue points, respectively. We notice that our measurements are fairly consistent with that from Johnston et al. (2021), shown as orange curves. The error bars are the standard deviation derived from the covariance matrix given by the OneCovariance code with the same redshift distribution and the best-fit $\Omega_{\mathrm{m}}$ and galaxy biases to be determined below. The shaded regions are angular scales corresponding to physical scales smaller than $8h^{-1}\mathrm{Mpc}$ estimated at the mean redshift. We then fit $\{\Omega_{\mathrm{m}},b_{1},b_{2}\}$ in the linear model with the 2PCFs in the linear scale. We note that we don’t know the parameters a priori to compute the covariance matrix, so we do an iterative parameter fit. We choose the initial value of $\Omega_{\mathrm{m}}\,$=0.33 and $b_{1}=1.1$, $b_{2}=1.25$ to compute the covariance matrix. The pilot galaxy bias values are taken from van Uitert et al. (2018), which apply to a GAMA-like subsample of KV-450 with a similar redshift distribution to the bright sample used here. We define a Gaussian likelihood and run an MCMC to obtain the best-fit parameters, and then update the covariance matrix with them. We then run the MCMC again with the updated covariance matrix to obtain the posterior.
@@
The posteriors of the UR and OR cases are shown as pink and blue contours in Fig. 27. The mean value of the parameters in the converged MCMC chains and the $1-\sigma$ levels are summarized in Table 4. The theoretical $w(\theta)$ calculated from the best-fit parameters of the OR case is shown as dashed black curves in Fig. 26. The reduced $\chi^{2}$ value between the OR 2PCF and the best-fit 2PCF is 1.06, corresponding to a PTE of 0.38, indicating a good fit between the model and the data. All three parameters are constrained. Notably, the matter density is constrained in agreement with previous cosmological probes. The galaxy biases are constrained to be close to 1 with a slightly increasing trend. However, we note that the bias parameters are highly degenerate with $\sigma_{8}$, which is fixed in our case. To break this degeneracy we need to introduce matter field tracers such as cosmic shear or use the halo model like in Dvornik et al. (2023). This is the content of the ongoing KiDS-Legacy $3\times 2$pt and $6\times 2$pt projects.
@@
Table 4: Parameter fit from KiDS-1000 bright galaxy 2PCF.
@@
 [^22] The best-fit values are those with maximum posterior and the errors are the 68.3% credible level.
Figure 27: Posterior of $\Omega_{\mathrm{m}},b_{1},b_{2}$ fit by the 2PCF from KiDS-1000 bright samples. The deep shaded and lightly shaded contour is the 68% and 95% credible levels, respectively. The blue and pink contours correspond to the posterior from 2PCF corrected with OR and UR respectively.